# Meta Co-Training: Two Views are Better than One

### v2

@@ 5 Experiments
MCT (MLP) 99.6 94.8 40.1
Table 7: Additional Comparisons to REACT [52]. The authors include experiments for zero-shot performance and 10% of available labels. Results above are for 10% of labels. For Flowers102 this is 1-shot performance.
@@
MCT (MLP) 76.0 (58.1) 94.8 (91.7)
Table 8: Additional Comparisons to Semi-ViT using 10% (1%) of the available labels for training. For the iNaturalist task we use only the 1010 most frequent classes as in [17].
Additional Datasets.   In Tables 7 and 8 we compare meta co-training and co-training to existing SotA approaches that leverage unlabeled data directly during training (e.g., not just part of a pretext task). These experiments support our hypothesis that the primary cause of the failure of co-training is the insufficiency of one or more views. The closer to equivalent in performance individual views are, the less the performance suffers. When the two views perform similarly and well, the performance improves. Section A in the appendix provides more information on these additional datasets; Figure 9 provides an example where co-training performs as expected. Meta co-training is in general more robust to view imbalance and provides better results.
@@ 6 Conclusion and Future Work
We presented a novel method of view construction for co-training approaches based on pretext tasks. We showed that the views themselves likely contain independent information and that models trained on these views can complement each other to produce stronger classifiers. However, these views appear to be insufficient for co-training on the ImageNet task. We proposed meta co-training, a novel algorithm for co-training to address the challenges faced by the co-training algorithm. We showed that our algorithm outperforms co-training in general, and establishes a new SotA top-1 classification accuracy on the ImageNet-10% dataset as well as new SotA few-shot top-1 classification accuracy on several other fine-grained image classification datasets.
For future work we expect that fine-tuning the representation learners that we used for co-training and meta co-training would yield better results; which should be in line to observations from papers proposing foundation models that report performance benefits of fine-tuning those models on downstream tasks. Furthermore, we did not make use of the large available unlabeled datasets for computer vision in this study. We suspect that giving our models access to a large repository of unlabeled data as well as the entirety of the labeled data for any dataset on which we performed experiments would result in a performance boost. Finally, our list of pretext tasks was not exhaustive. It is an interesting question for future work which other tasks and models might yield useful views.
Acknowledgements.   This material is based upon work supported by the National Science Foundation under Grant No. ICER-2019758. This work is part of the NSF AI Institute for Research on Trustworthy AI in Weather, Climate, and Coastal Oceanography (AI2ES).
@@
Appendix A Further Analysis of Meta Co-Training and Co-Training
Figures 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, and 15 include the top-1 accuracy of co-training using the best (CLIP and DINO) views found for the ImageNet task. In all cases except Figure 9, co-training reaches its maximum performance in the first iteration. Note that there is no 1% run for Flowers102 because there are only 10 examples for each class in the training set (see the discussion in Section B).
A.1 Further Analysis of Meta Co-Training Experiments on ImageNet
In Figures 5 and 6 are shown aggregate results comparable to those in Figures 1 and 2 in the main text, but for meta co-training. The min, max, and mean performance of any of the ten pairs is shown. Note that these are not three runs, but aggregate statistics over all runs, so the minimum and maximum points on the graph for one step may not correspond to the same pair as another step.
Figure 5: Aggregate statistics of meta co-training across all views. Models are trained on 10% of the ImageNet labels. The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
Figure 6: Aggregate statistics of meta co-training across all views. Models are trained on 1% of the ImageNet labels. The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
In Figures 7 and 8 are shown results comparable to those in Figures 3 and  4 in the main text, but for meta co-training. Unlike co-training, meta co-training performance increases after the supervised warmup.
Figure 7: Meta co-training using the CLIP and DINOv2 views as a function of the training step. Models are trained on 10% of the ImageNet labels.
Figure 8: Meta co-training using the CLIP and DINOv2 views as a function of the training step. Models are trained on 1% of the ImageNet labels.
A.2 Co-Training and Meta Co-Training on Food101
In Figures 9 and  10 we show the top-1 accuracy of the iterations of co-training on the Food101 classification task. Co-training performs as expected with the pseudo labels allowing the other model to improve in the 10% case, but in the 1% case it appears that one or more of the views was insufficient and the performance decreases rapidly after the first iteration. Co-training achieves top-1 accuracy of 94.7% (resp., 83.9%) when 10% (resp., 1%) of the available labels for training are used from the Food101 dataset.
Figure 9: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available Food101 labels are used for training. The method exhibits performance improvement over multiple iterations of pseudo-labeling and retraining.
Figure 10: The top-1 accuracy of the co-training predictions for each iteration of co-training. 1% of available Food101 labels are used for training.
Similarly, meta co-training achieves top-1 accuracy of 94.8% (resp., 91.7%) when 10% (resp., 1%) of the available training examples of the Food101 dataset are used.
These results are compared against each other as well as against REACT and Semi-ViT in Tables 7 and 8 respectively.
A.3 Co-Training and Meta Co-Training on iNaturalist
In Figures 11 and 12 we show the top-1 accuracy of the iterations of co-training on the iNaturalist dataset. In both splits we see a steady decline in top-1 accuracy after the first iteration. Co-training achieves top-1 accuracy of 59.7% (resp., 29.5%) when only 10% (resp., 1%) of the training examples of the iNaturalist dataset are used.
Figure 11: The top-1 accuracy of the co-training predictions for each iteration of co-training. Only the 1010 most frequent classes are included. 10% of available iNaturalist labels from those classes are used for training.
Figure 12: The top-1 accuracy of the co-training predictions for each iteration of co-training. Only the 1010 most frequent classes are included. 1% of available iNaturalist labels from those classes are used for training.
Along these lines meta co-training achieves 76.0% (resp., 58.1%) when only 10% (resp., 1%) of the training examples of the iNaturalist dataset are used.
These results are compared against each other as well as against REACT and Semi-ViT in Tables 7 and 8 respectively.
A.4 Co-Training and Meta Co-Training on CTFGVCAircraft
In Figures 13 and 14 we show the top-1 accuracy of the iterations of co-training on the CTFGVCAircraft dataset. In both cases both views have low accuracy, so it is unsurprising that in subsequent iterations the accuracy decreases rapidly. In particular, co-training achieves top-1 accuracy of 36.4% (resp., 21.6%) when only 10% (resp., 1%) of the training examples of the CTFGVCAircraft dataset are used.
Figure 13: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available FGVCAircraft labels are used for training.
Figure 14: The top-1 accuracy of the co-training predictions for each iteration of co-training. 1% of available FGVCAircraft labels are used for training.
Along these lines meta co-training achieves 40.1% (resp., 21.5%) when only 10% (resp., 1%) of the training examples of the CTFGVCAircraft dataset are used.
The results corresponding to the 10% case of co-training and meta co-training are compared against REACT in Table 7, since the authors of [52] report results on zero-shot and 10% cases only. However, for completeness, we have decided to include Figure 14 which corresponds to the experiments that we did using only 1% of the labeled training data available for the CTFGVCAircraft dataset.
Dataset #train #test #classes
FGVCAircraft $3$ $3$ 100
Flowers102 1020 6149 102
Food101 75,750 25,250 101
iNaturalist 175,489 29,083 1010
ImageNet 1,281,167 50,000 1000
Table 9: Sizes of datasets used. Subsets drawn for training are always drawn in a stratified way to maintain the class distribution of the original data. For ImageNet we use the split published with the SimCLRv2 [24] repository.
config MCT Co-Training
optimizer Adam Adam
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.999$ $\beta_{1},\beta_{2}=0.9,0.999$
base learning rate 1e-4 1e-4
batch size 4096 4096
learning rate schedule Reduce On Plateau Reduce On Plateau
schedule patience 10 steps 10 steps
decay factor 0.5 0.5
training steps 1000 200 (per iteration)
warmup steps 200 n/a
Table 10: Hyperparmeter configuration for meta co-training and co-training. Other models used (MLP, linear) are trained for the same number of steps as meta co-training, with the same optimization parameters.
A.5 Co-Training and Meta Co-Training on Flowers102
As explained earlier, the Flowers102 dataset has only 10 training examples for each class and hence 10% corresponds to 1-shot learning. In this context, in Figure 15 we give the performance of the iterations of co-training on the Flowers102 dataset. Even though the DINO view performs very well in the first iteration, there is a large gap between the performance of the CLIP view and the DINO view which leads to decreased performance in subsequent iterations. Co-training achieves top-1 accuracy of 99.2% when only 10% of the training examples of the Flowers102 dataset are used for training (i.e., a single labeled image per class).
In the same context of 1-shot learning, meta co-training achieves top-1 accuracy of 99.6%.
Both of the above results are compared against each other and against REACT in Table 7.
Figure 15: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available Flowers102 labels are used for training.
Appendix B Further Information on the Datasets and the Splits
Table 9 includes a description of the different sizes of the datasets used for training and testing, as well as the number of cases for each dataset. Note that in the 10% experiments on the Flowers102 dataset each class has only 1 label. All datasets are approximately class-balanced with the exception of the iNaturalist dataset. In all cases we maintain the original class distribution when sampling subsets. All subsets are created with seeded randomness of a common seed (13) which ensures a fair comparison between methods.
Appendix C Hyperparameter Configuration
@@
The relevant hyperparameters for training are included in Table 10 for co-training and meta co-training (MCT).

### v5

@@ 4.3 Experiments on Additional Datasets
Meta Co-Training (MLP) 99.6 94.8 40.1
Table 10: Additional comparisons to Semi-ViT using 10% (1%) of the available labels for training. For the iNaturalist task we use only the 1010 most frequent classes.
@@
Meta Co-Training (MLP) 76.0 (58.1) 94.8 (91.7)
Figure 4: The top-1 accuracy of the CT predictions for each iteration of CT. 10% of available Food101 labels are used for training. The method exhibits performance improvement over multiple iterations of pseudo-labeling and retraining.
5 Related Work
While SSL has become very relevant in recent years with the influx of the data outpacing the availability of human labels, the paradigm and its usefulness have been studied at least since the 1960’s; e.g., [25, 16, 48, 22]. Among the main approaches for SSL that are not directly related to our work, one popular direction encourages entropy minimization and consistency regularization using data augmentation such as MixMatch [4], UDA [51], ReMixMatch [3], FixMatch [49], and FlexMatch [53].
Self-Training and Student-Teacher Frameworks.
One of the earliest ideas for SSL is that of self-training [48, 22]. This idea is still very popular with different student-teacher frameworks [55, 52, 43, 35, 24, 9]. This technique is also frequently referred to as pseudo-labeling [33] as the student tries to improve its performance by making predictions on the unlabeled data, then integrating the most confident predictions to the labeled set, and subsequently retraining using the broader labeled set. In addition, one can combine ideas; e.g., self-training with clustering [30], or create ensembles via self-training [31], potentially using graph-based information [38]. An issue with these approaches is that of confirmation bias [1], where the student cannot improve a lot due to incorrect labeling that occurs either because of self-training, or due to a fallible teacher.
Co-training Algorithms.
Co-training was introduced by Blum and Mitchell [5] and their paradigmatic example was classifying web pages using two different views: (i) the bags of words based on the content that the webpages had, and (ii) the bag of words formed by the words used in hyperlinks pointing to these webpages. In our context the two different views are obtained via different embeddings that we obtain when we use images as inputs to different foundation models and observe the resulting embeddings.
Co-training has been very successful with wide applicability [20] and can outperform learning algorithms that use only single-view data [40]. Furthermore, the generalization ability of co-training is well-motivated theoretically [5, 15, 2, 14]. There are several natural extensions of the base idea, such as methods for constructing different views [50, 7, 11, 8], including connections to active learning [19], methods that are specific to deep learning [44], and other extensions that exploit more than two views on data; e.g., [54].
@@ 6 Conclusion and Future Work
We proposed MCT, a novel semi-supervised learning algorithm. We showed that our algorithm outperforms co-training in general and establishes a new SoTA top-1 classification accuracy on the ImageNet 10%. We also tested the performance of deep ensembles on ImageNet 10% and 1%. In both cases MCT achieved better results, demonstrating that MCT provides utility beyond what can be attributed to ensembling. In addition, using MCT, we were able to achieve new SoTA few-shot top-1 classification accuracy on other popular CV benchmark datasets, including one-shot top-1 classification accuracy for the Flowers102 dataset. Furthermore, we presented evidence that self-supervised learned representations are good candidates for views. We showed that the views constructed this way likely contain independent information and that models trained on these views, or concatenation of these views, can complement each other to produce stronger classifiers.
Our experiments indicate that our method performs best when it has access to enough labeled data to produce strong views for pseudo-labeling. For some datasets this may preclude its application in one-shot settings. Further, if no sufficient and independent views are available and none can be constructed then we do not expect our model to perform meaningfully better than MPL.
For future work we expect that fine-tuning the representation learners that we used for CT and MCT would improve results. Papers proposing foundation models observe performance benefits when fine-tuning those models on downstream tasks. Furthermore, we did not make use of the large available unlabeled datasets for CV in this study. Giving MCT access to a large repository of unlabeled data, as well as the entirety of the labeled data, should result in a performance boost. Finally, our list of pretext tasks was not exhaustive. It is an interesting question for future work which other tasks and models might yield useful views.
7 Acknowledgments
This material is based upon work supported by the U.S. National Science Foundation under Grant No. RISE-2019758. This work is part of the NSF AI Institute for Research on Trustworthy AI in Weather, Climate, and Coastal Oceanography (NSF AI2ES).
@@ 7 Acknowledgments
Appendix A Further Details on Related Methods
Our work is inspired by two SSL methods: meta pseudo labels (MPL) [43] and co-training [5]. We provide more information on these methods below. This discussion should provide further insight for Algorithm 1.
Meta Pseudo Labels.
The hope is that updated student parameters $\theta_{S}^{\text{PL}}$ will behave well on the labeled data $L=(X_{L},Y_{L})=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$; that is, the loss $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)=\ell\left(Y_{L},f_{\theta_{S}^{\text{PL}}}(X_{L})\right)$ is small, where $\mathcal{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ is defined so that the dependence between $\theta_{S}$ and $\theta_{T}$ is explicit.
How we might exploit the dependency between $\theta_{T}$ and $\theta_{S}$ might not be immediately obvious. Eventually the models are optimized using $\theta_{S}^{\text{PL}}\approx\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)$ as a practical approximation [36, 21] which leads to the tractable teacher objective:
For a more detailed explanation of this relationship we refer the interested reader to [43].
Co-Training.
Co-training relies on the existence of two different views of the instance space $\mathcal{X}$; i.e., $\mathcal{X}=\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$. Hence, an instance $x\in\mathcal{X}$ has the form $x=(x^{(1)},x^{(2)})\in\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$. We call the views $x^{(1)}$ and $x^{(2)}$ complementary. Co-training requires two assumptions in order to work. First, it is assumed that, provided enough many examples, each separate view is enough for learning a meaningful model. The second is the conditional independence assumption [5], which states that $x^{(1)}$ and $x^{(2)}$ are conditionally independent given the label of the instance $x=(x^{(1)},x^{(2)})$. Thus, with co-training, two models $f^{(1)}$ and $f^{(2)}$ are learnt, which help each other using different information captured from the different views of each instance. Below we provide more information.
Given a dataset $S=L\cup U$ (Section 2 has notation), we can have access to two different views of the dataset due to the natural partitioning of the instance space $\mathcal{X}$; that is, $S^{(1)}=L^{(1)}\cup U^{(1)}$ and $S^{(2)}=L^{(2)}\cup U^{(2)}$. Initially two models $f^{(1)}$ and $f^{(2)}$ are learnt using a supervised learning algorithm based on the labeled sets $L^{(1)}$ and $L^{(2)}$ respectively. Learning proceeds in iterations, where in each iteration a number of unlabeled instances are integrated with the labeled set by assigning pseudo-labels to them that are predicted by $f^{(1)}$ and $f^{(2)}$, so that improved models $f^{(1)}$ and $f^{(2)}$ can be learnt by retraining on the augmented labeled sets. In particular, in each iteration, each model $f^{(v)}$ predicts a label for each view $x^{(v)}\in U^{(v)}$. The top-$k$ most confident predictions by model $f^{(v)}$ are used to provide a pseudo-label for the respective complementary instances and then these pseudo-labeled examples are integrated in to the labeled set for the complementary view. In the next iteration a new model is trained based on the augmented labeled set. Both views of the used instances are dropped from $U$ even if only one of the views may be used for augmenting the appropriate labeled set. This process repeats until the unlabeled sets $U^{(v)}$ are exhausted, yielding models $f^{(1)}$ and $f^{(2)}$ corresponding to the two views. Co-Training is evaluated using the joint prediction of these models which is the re-normalized element-wise product of the two predictions:
where $\otimes$ is the Hadamard (element-wise) product of two vectors; that is, for vectors $\alpha,\beta\in\mathbb{R}^{d}$ their Hadamard (element-wise) product is the vector $\gamma=\alpha\otimes\beta$, where $\gamma_{j}=\alpha_{j}\cdot\beta_{j}$ for $j\in\{1,\ldots,d\}$.
Appendix B Hyperparameter Configuration
@@
The relevant hyperparameters for training are included in Table 11 for co-training (CT), meta co-training (MCT), and deep ensembles.
Table 11: Hyperparmeter configuration for CT, MCT, and deep ensemble. Other models used (MLP, linear) are trained for the same number of steps as MCT, with the same optimization parameters.
config Co-Training Meta Co-Training Deep Ensemble
optimizer Adam Adam Adam
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.999$ $\beta_{1},\beta_{2}=0.9,0.999$ $\beta_{1},\beta_{2}=0.9,0.999$
base learning rate 1e-4 1e-4 1e-4
batch size 4096 4096 4096
learning rate schedule Reduce On Plateau Reduce On Plateau Reduce On Plateau
schedule patience 10 steps 10 steps 10 steps
decay factor 0.5 0.5 0.5
training steps 200 (per iteration) 1000 1000 (per model)
warmup steps n/a 200 n/a
Appendix C Omitted Experimental Analysis on ImageNet
In this section we provide some figures illustrating the experimental results on ImageNet 10% and ImageNet 1% that we did not include into the main text due to space limitations.
In Figures 5 and 6 are shown results comparable to those in Figures 2 and 3 but for the ImageNet-1% dataset. As a reminder these figures are showing top-1 accuracy for CT and MCT when the constituent views used in the embeddings are CLIP and DINOv2. Similar to the case of ImageNet 10% we see that the CT performance decreases after the supervised warmup. Unlike CT, MCT performance increases after the supervised warmup.
Figure 5: Top-1 accuracy of CT iterations on the CLIP and DINOv2 views for the ImageNet 1% dataset.
Figure 6: MCT using the CLIP and DINOv2 views as a function of the training step. Models are trained on 1% of the ImageNet labels.
Figure 7: Top-1 accuracy of CT iterations for the ImageNet 10% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
Figure 8: Aggregate statistics of MCT across all views. Models are trained on 10% of the ImageNet labels. The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
Figure 9: Top-1 accuracy of CT iterations for the ImageNet 1% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
Figure 10: Aggregate statistics of MCT across all views. Models are trained on 1% of the ImageNet labels. The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
In Figures 7 and 8 we show the min, max, and mean performance of all of the ten pairs of views for the ImageNet-10% dataset. Note that these are not three runs, but aggregate statistics over all runs, so the minimum and maximum points on the graph for one step may not correspond to the same pair as another step. Figures 9 and 10 are comparable to Figures 7 and 8, but for the ImageNet-1% dataset.
Appendix D Further Details on Co-Training Iterations on Additional Datasets
In this section we provide further details on the performance of pure CT on additional datasets, using the two views that appear to have the highest information content among those that we tried; i.e., CLIP and DINOv2. These experiments support our hypothesis that the primary cause of the failure of CT is that one or more views can be weak and do not suffice to learn a model that achieves reasonably high accuracy. The closer to equivalent in performance individual views are, the less the performance suffers. When the two views perform similarly and well, the performance improves.
Figure 11 and Figure 12 show CT iteration performance on the FGVC Aircraft dataset using 10% and 1% of labels respectively.
Figure 11: The top-1 accuracy of the CT predictions for each iteration of CT. 10% of available FGVC Aircraft labels are used for training. The method exhibits performance improvement over multiple iterations of pseudo-labeling and retraining.
Figure 12: The top-1 accuracy of the CT predictions for each iteration of CT. 1% of available FGVC Aircraft labels are used for training.
Neither view performs well and there is a large difference between the performance of each view. The performance of CT quickly collapses.
Figure 4, which was already presented in the main body of this work, provides an example where CT performs as expected and accuracy increases as CT progresses, though it can be seen from Figure 13 that the degeneration still occurs in the case with less labels.
In Figure 13 we show the iterations of CT for Food101-1%. Figure 4 with the iterations for Food101-10% appears in 4. The 1-shot results for Flowers-102 (10%) appear in Figure 14. The iNaturalist-2017 results for 10% and 1% respectively appear in Figures 15, 16.
Figure 13: The top-1 accuracy of the CT predictions for each iteration of CT. 1% of available Food101 labels are used for training.
Figure 14: The top-1 accuracy of the CT predictions for each iteration of CT. 10% of available Flowers102 labels are used for training.
Figure 15: The top-1 accuracy of the CT predictions for each iteration of CT. Only the 1010 most frequent classes are included. 10% of available iNaturalist labels from those classes are used for training.
Figure 16: The top-1 accuracy of the CT predictions for each iteration of CT. Only the 1010 most frequent classes are included. 1% of available iNaturalist labels from those classes are used for training.
Main Takeaway.
The main takeaway of these additional experiments, apart from establishing new SoTA results in datasets beyond ImageNet, is that MCT is in general more robust to view imbalance and provides better results than CT.